Kates age 5 years ago, when doubled, is equal to triple her age 10 years ago
Answer by macston(5194)

K=Kate's age
2(K-5yrs)=3(K-10yrs)
2K-10yrs=3K-30yrs Add 30 yrs to each side.
2K+20yrs=3K Subtract 2K from each side.
20yrs=K ANSWER: Kate is 20 years old.
CHECK:
2(K-5yrs)=3(K-10yrs)
2(20yrs-5yrs)=3(20yrs-10yrs)
2(15yrs)=3(10yrs)
30yrs=30yrs